# Moonshot: Optimizing Chain-Based Rotating Leader BFT via Optimistic Proposals

### v1

@@ Appendix C Commit Moonshot Security Analysis
As in the base case, by the direct commit rule, $P_{i}$ must have received $12f+$ 1 $\langle\mathsf{commit},H(B_{k}),v\rangle_{*}$, at least $f+1$ of which must have been sent by a set $H_{2}$ of distinct honest nodes. Hence, by the pre-commit rule, $H_{2}$ must have received $\mathcal{C}_{v}(B_{k})$ whilst having ${\sf timeout\_view}_{i}<v$. Thus, by 1, every $\mathsf{timeout}_{v^{*}}$ message sent by $H_{2}$ will necessarily contain $\mathcal{C}_{v}(B_{k})$ or higher. Therefore, because $v^{\prime}-1\geq v+1$ and since by 2 $H_{1}$ and $H_{2}$ must have at least one node in common, the highest ranked block certificate of $\mathcal{TC}_{v^{\prime}-1}$ must have a rank of at least as great as $\mathcal{C}_{v}(B_{k})$. Thus, $v^{\prime\prime}\geq v$. Therefore, since $v\leq v^{\prime\prime}<v^{\prime}$, by the induction hypothesis, $B_{k^{\prime\prime}}$ extends $B_{k}$, so $B_{k^{\prime}}$ also extends $B_{k}$. ∎

### v2

@@ Appendix C Commit Moonshot Security Analysis
By the alternative direct commit rule, $P_{i}$ must have received $12f+$ 1 distinct $\langle\mathsf{commit},H(B_{k}),v\rangle_{*}$, at least $f+1$ of which must have been sent by a set $H_{2}$ of distinct honest nodes. Therefore, by 13, every timeout certificate for a view greater than $v$ must contain a block certificate for $v$ or higher. Thus, because $v^{\prime}-1\geq v+1$ and since by 2 $H_{1}$ and $H_{2}$ must have at least one node in common, the highest ranked block certificate of $\mathcal{TC}_{v^{\prime}-1}$ must have a rank of at least as great as $\mathcal{C}_{v}(B_{k})$. Hence, $v^{\prime\prime}\geq v$. Therefore, since $v\leq v^{\prime\prime}<v^{\prime}$, by the induction hypothesis, $B_{k^{\prime\prime}}$ extends $B_{k}$, so $B_{k^{\prime}}$ also extends $B_{k}$. ∎
Claim 14 (Single Leader Commit).
If the first honest node to enter view $v$ does so at time $t$ after GST and $L_{v}$ is honest, then all honest nodes commit one of its proposals before $t+4\Delta$.
Proof.
By Lemma 12, all honest nodes receive $\mathcal{C}_{v}(B_{k})$ for some block $B_{k}$ proposed by $L_{v}$, before $t+3\Delta$. Therefore, by the alternative direct commit rule, if they all multicast $\langle\mathsf{commit},H(B_{k}),v\rangle_{*}$ upon receiving this certificate, then all honest nodes will commit $B_{k}$ before $t+4\Delta$. Otherwise, at least one honest node, say $P_{i}$, must fail to send $\langle\mathsf{commit},H(B_{k}),v\rangle_{*}$ before $t+3\Delta$. However, by 10, no honest node can have its view timer expire before this time, so $P_{i}$ cannot have ${\sf timeout\_view}_{i}\geq v$ upon receiving $\mathcal{C}_{v}(B_{k})$. Thus, by the pre-commit rules, $P_{i}$ must neither be in view $v$ or lower, nor have multicasted a commit vote for any descendant of $B_{k}$, upon receiving $\mathcal{C}_{v}(B_{k})$. Let $v^{\prime}$ denote the view that $P_{i}$ was in upon receiving this certificate. Since, by Corollary 1, no timeout certificate can exist for $v$ or higher before $t+3\Delta$, $P_{i}$ must have entered $v^{\prime}$ via $\mathcal{C}_{v^{\prime}-1}(B_{l})$. Moreover, by the direct pre-commit rule, it would have multicasted $\langle\mathsf{commit},H(B_{l}),v^{\prime}-1\rangle_{j}$ upon receiving this certificate. Therefore, $B_{l}$ must not be a descendant of $B_{k}$. However, it must also be true that $v^{\prime}>v+1$ and that all block certificates for the views between $v^{\prime}$ and $v$ must be either $\mathcal{C}^{o}$ or $\mathcal{C}^{n}$. Therefore, by the corresponding vote rules, $B_{l}$ must be a descendant of $B_{k}$, contradicting the former conclusion that it must not be. Therefore, all honest nodes will multicast $\langle\mathsf{commit},H(B_{k}),v\rangle_{*}$ before $t+3\Delta$, so all honest nodes will commit $B_{k}$ before $t+4\Delta$. ∎
Appendix D Further Optimizing View Length
The view lengths of Pipelined Moonshot and Commit Moonshot can be further reduced in views without valid proposals by applying the modifications presented in Figure 10. These modifications leverage the fact that these protocols guarantee that all nodes will receive a valid proposal from an honest $L_{v}$ within $2\Delta$ of the first honest node entering $v$ after GST (see 11). Consequently, if a node has to wait more than $2\Delta$ to receive a valid proposal, it can be confident either that the network is asynchronous or that the leader is Byzantine and therefore has reason to begin the view change process. Similarly, if it votes for such a proposal then it should expect to construct a certificate for the included block within $3\Delta$ of the first honest node entering the view (see Lemma 12) and thus should increase its view timer by $\Delta$ upon voting to allow sufficient time for the votes of its peers to arrive. While this latter modification preserves liveness, it means that Pipelined Moonshot and Commit Moonshot will only exhibit a view length of $2\Delta$ in views in which no honest nodes vote for any block. Accordingly, it is trivial for Byzantine leaders to ensure that their views retain the original view length of $3\Delta$. Even so, this modification represents a meaningful optimization in the crash fault tolerant setting.
The $\tau$ of Pipelined Moonshot and Commit Moonshot can be further reduced by modifying the protocol for for $P_{i}$ presented in Figure 3 as follows: 1. Reset Timer. Upon entering $v$, reset ${\sf view\text{-}timer}_{i}$ to $2\Delta$ and start counting down. This replaces the corresponding logic given in the Advance View rule. 2. Increase Timer. Upon voting in $v$, increase ${\sf view\text{-}timer}_{i}$ by $\Delta$.
Figure 10: Further Optimizing View Length